OpenStudy (anonymous):

Solve the following algebraic proportion for x, in terms of a and b, showing the steps in your work. x+7/a = 3x+1/b when a = 5 and b = 3, what is the value of x?

OpenStudy (campbell_st):

so you have a problem \[\frac{x + 7}{5} = \frac{3x +1}{3}\] 1st step is to cross multiply to remove the denominators 3(x + 7) = 5(3x + 1) now you need to distribute then solve for x hope it helps

OpenStudy (anonymous):

which numbers to I cross multiply by first, and thank you for your help

OpenStudy (campbell_st):

well the cross multiplication as been done for you... above 3(x + 7) = 5(3x + 1) now distribute